OpenStudy (anonymous):

simplify the expression shown below ln(e^3 x^4)

OpenStudy (anonymous):

would i need to separate them soooo lne^3 / lnx^4 ????

OpenStudy (anonymous):

\[\ln(e^3x^4)\]

OpenStudy (anonymous):

or would it be lne^3 + lnx^4 ???

OpenStudy (anonymous):

correct carry on

OpenStudy (ikram002p):

Hint :- ln(A .B)=ln(A)+ln (B) so ur correct its lne^3 + lnx^4

OpenStudy (ikram002p):

also ln (A^n)=n ln (A )

OpenStudy (nincompoop):

that is still not solved you just did it in expanded form

OpenStudy (ikram002p):

"simplify the expression shown below " it said simplify not solve

OpenStudy (ikram002p):

so go with hint 2 ln (A^n)=n ln (A ) lne^3 + lnx^4 = 3 ln e + 4 ln x hint 3 ln e =1

OpenStudy (anonymous):

sooooooo... 3+4lnx

OpenStudy (ikram002p):

yep :D

OpenStudy (anonymous):

yay soo that as simplify as it gets ? @ikram002p

OpenStudy (ikram002p):

well yeah thats what i got ,if u have another method/solution i'd be glad to know\learn :D

OpenStudy (anonymous):

ahahahhaa i dont im barely hanging on with logs lol thank you n_n @ikram002p

OpenStudy (ikram002p):

np :)